OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 15x - y = -6 5x - 3y = -13 (1 point) 5(-15x - 6) - 3y = -13 5(15x + 6) - 3y = -13 5x - 3(15x + 6) = -13 5x - 3(-15x - 6) = -13

OpenStudy (anonymous):

we learned this last quarter but it wont let me go back to the class on my online school it said session has expired, i know ho to solve for x and y in systems of equations i dont know how to use the substitution mathod

OpenStudy (anonymous):

how to* method*

OpenStudy (ranga):

From the first equation find y in terms of y first: 15x - y = -6 y = 15x + 6 Put this y in the second equation: 5x - 3y = -13 5x - 3(15x + 6) = -13 5x - 45x - 18 = -13 -40x = 18 - 13 = 5 x = -5/40 = -1/8 put it in y = 15x + 6 y = 15(-1/8) + 6 y = (-15 + 48)/8 = 33/8

OpenStudy (ranga):

find y in terms of *x*

OpenStudy (anonymous):

1 sec im looking over it im just slow i have to read it twice

OpenStudy (ranga):

Oh, they don't want the solution, just the substitution. So you can stop at line 5: 5x - 3(15x + 6) = -13

OpenStudy (anonymous):

ok, hold on 1 sec (:

OpenStudy (anonymous):

im still confused its not clicking in ym head

OpenStudy (anonymous):

my*

OpenStudy (anonymous):

i dont understand how i come to a conclusion of one of those answers.

OpenStudy (anonymous):

wait oh o.e

OpenStudy (anonymous):

i was looking at the wrong problem

OpenStudy (anonymous):

nvm i get it now XD

OpenStudy (ranga):

ok.

OpenStudy (anonymous):

i was looking at 4. Solve the following system of equations. 2x + 4y - 3z = -7 3x + y + 4z = -12 x + 3y + 4z = 4 (1 point) (-6, -2, -1) (6, -2, -1) (-6, 2, 1) (6, 2, 1)

OpenStudy (anonymous):

ty for your help ^.^